OpenStudy (anonymous):

(a+8)/(a-1) +(a+4)/(a+1) - (8a+2)/(a^2-1)

OpenStudy (anonymous):

is this all one equation?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I have to simplify it.

OpenStudy (anonymous):

I don't need to know a.

OpenStudy (anonymous):

i guess would you add all the three equations up and then divide them?

OpenStudy (anonymous):

so it would be (2a+12)/(2a+5)-(8a+2)/(a*2-1) and then you simplify that?

OpenStudy (jack1):

@L-Lawliet-L to add these fractions, you need to have the same base (denominator)

OpenStudy (jack1):

so your bases are: a-1 a+1 and a^2 - 1 so convert them to the same base (use a^2 - 1)

OpenStudy (jack1):

(a+8)/(a-1) + (a+4)/(a+1) - (8a+2)/(a^2-1) so first fraction; multiply top and bottom by (a+1): (a+8) * (a+1) = (a^2 + 8a + 1a + 8) ----- ------------------ (a-1) * (a+1) = a^2 - 1

OpenStudy (jack1):

now 2nd fraction; multiply top and bottom by (a-1): (a+4) * (a-1) = ( ...??? @L-Lawliet-L ) ------ ------------------ (a+1) * (a-1) = a^2 - 1

OpenStudy (jack1):

anyway, now all 3 have the same denominator, so you can add the numerators across directly (so just all all like terms on the top lines)